INFRARED SPECTRA OF BENZENE-DIACETYLENE CLUSTERS: C_{6}H_{6} · C_{4}H_{2} and C_{6}H_{6} · (C_{4}H_{2})_{2}

\underline{CHRISTOPHER RAMOS}, PAUL R. WINTER, and TIMOTHY S. ZWIER, Department of Chemistry, Purdue University, West Lafayette, IN 47907-1393.

Resonant two photon ionization spectra of the benzene:diacetylene 1:1 cluster have been taken across the 6^{1}_{0} transition region of the cluster with time-of-flight mass spectrometry detection. IR/UV holeburning has shown that one of the peaks in the spectrum is the benzene:diacetylene 1:2 cluster. The resonant ion dip infrared spectra of both the 1:1 cluster and the 1:2 cluster have been recorded. Possible structures of these clusters will be discussed.